# Either a Confidence Interval Covers, or It Doesn’t (Or Does It?): A Model-Based View of Ex-Post Coverage Probability

### v1

@@
In Neyman’s original formulation, a $(1-\alpha)$ confidence interval procedure is justified by its long-run coverage properties, and a single realized interval is to be described only by the slogan that it either covers the parameter or it does not. On this view, post-data probability statements about the coverage of an individual interval are taken to be conceptually out of bounds. In this paper, I present two kinds of arguments against treating that “either-or” reading as the only legitimate interpretation of confidence. The first is informal, via a set of thought experiments in which the same joint probability model is used to compute both forward-looking and backward-looking probabilities for occurred-but-unobserved events. The second is more formal, recasting the standard confidence-interval construction in terms of infinite sequences of trials and their associated $0/1$ coverage indicators. In that representation, the design-level coverage probability $1-\alpha$ and the degenerate conditional probabilities given the full data appear simply as different conditioning levels of the same model. I argue that a strict behavioristic reading that privileges only the latter is in tension with the very mathematical machinery used to define long-run error rates. I then sketch an alternative view of confidence as a predictive probability (or forecast) about the coverage indicator, together with a simple normative rule for when intermediate probabilities for single coverage events should be allowed.
@@ 1.1 Background
The tension between the accepted interpretation of CIs and the supposedly-fallacious one can be recast more generally as a statement about events we know have occurred, but whose outcomes we have not observed [25]: for frequentists, randomness, and thus probability, lives in the sampling process and not our knowledge of the outcomes, and so once a sample has been drawn, the ex-ante (i.e., ”pre-data”) probability has collapsed to some value in $\{0,1\}$. Although this is mathematically true, it should also ring alarm bells for practicing statisticians (but, in the case of CIs, never seems to), because we happily use frequentist methods for statistical inference in exactly this kind of real-world scenario. Take, for example, the case of medical diagnosis: given that a patient tests positive on, say, a rapid diagnostic test for the influenza virus, what is the probability that she actually has it? If we stay consistent with our interpretation of CIs, we should also say no probability statement may now be made: given her true underlying health state, the patient either does or does not have the flu, and there is no probability left to assign, because all of the randomness in the sampling process has now been exhausted. Clearly, though, following the ”either-or” logic here would ruin the clinical value of the diagnostic test in guiding care, and it would obviate the effort epidemiologists and statisticians put in to estimating the test’s positive predictive value ($PPV$) in the first place, neither of which seems particularly desirable.
@@ 1.2 Paper overview and contributions
In what follows, I deliberately adopt a rather strict reading of Neyman’s slogan–in a nutshell, that ex post probabilities of coverage are entirely out of bounds–and treat it as a normative rule (to be fair, though, this is how many, if not most, instructive pieces handle the interpretation, see e.g., [2, 22, 29, 10, 1, 18] for examples). The arguments will form a kind of reductio: if we really insist on this kind of rule, we run into uncomfortable, if not untenable, constraints on other frequentist uses of probability. The alternative I suggest, and perhaps this paper’s main contribution, is that we keep Neyman’s (very useful) conception of long-run error control, but that we loosen what we are allowed to say about individual coverage events, especially ex post.
@@ 3.1 Confidence intervals and coverage indicators
with $Z_{i}(\omega)\in\{0,1\}$ for every $i$. From this viewpoint, a single realized world $\omega$ contains a fully-fixed infinite sequence of intervals and coverage outcomes; the only randomness is in which index $i$ of this sequence is consulted when the procedure is run. The strong law of large numbers applied to $(Z_{i})_{i\geq 1}$ yields
@@ 4.2 What is confidence, then?
The discussion above suggests that part of the trouble comes from conflating distinct layers of probability that the model makes available. We clearly have at least two model-based layers to work with when interpreting confidence intervals: at the design level, there is the unconditional coverage probability $1-\alpha$ attached to the procedure as a whole; and at the ex post level, conditioned on the truth value of the coverage event for a particular interval, there are the degenerate probabilities 0 and $1$ corresponding to non-coverage and coverage, respectively. Both are properties of the same probability measure $P_{\theta}$, viewed at different conditioning levels. The design-level quantity arguably holds up better under scrutiny, as it avoids the multi collective complications noted above. For some distributions and estimators, though, a bare report of $1-\alpha$ can feel somewhat strange ex post. The “trivial” interval, which returns the real line with probability $1-\alpha$ and the empty set with probability $\alpha$, is perhaps the most familiar example, but Welch’s uniform confidence interval [31] and Basu’s construction based on a $\theta$-free $X$ [3] are also instructive.
@@
In such cases, I would suggest that, while the two model-based layers are certainly still in play–considered only as a random draw, any interval generated by the procedure does, by definition, inherit the design-level coverage probability—it may be more natural to admit a third layer of probability: that of a single-case predictive probability, which can be scored against the underlying $0/1$ coverage outcome. From this perspective, our $1-\alpha$ report ex ante is a forecast about the coverage indicator for the interval to be constructed, distinct from the interval’s realized coverage status under the model, and chosen so that it minimizes expected loss under a proper scoring rule over repeated sampling [9]. In a prequential reading [4], repeated use of the procedure produces a sequence of forecasts and outcomes
$(p_{1},Z_{1}),(p_{2},Z_{2}),\ldots,$
where $Z_{i}$ is the coverage indicator for the $i$th interval and $p_{i}$ is the predictive probability assigned to $Z_{i}=1$ given whatever information we choose to condition on. The forecast sequence $(p_{i})$ can then be evaluated by its long-run scores and calibration properties, independently of whether we identify $p_{i}$ with any particular $P_{\theta}$-probability. While our natural ex ante forecast would simply be $p_{i}=1-\alpha$, reflecting only the design, we might legitimately allow our ex post forecast to change if, as in the case of Welch’s uniform CI, features of both the design and the numerical value of the realized interval provide additional clues about coverage. In that richer information state, the predictive probability $p_{i}(X_{i})$ might not equal $1-\alpha$, but it would not also necessarily collapse to $\{0,1\}$: it would be a genuine intermediate forecast about the coverage event $Z_{i}$, subject to empirical assessment across repeated uses of the procedure.
@@
Along these lines, I suggest that this kind of forecast is, in essence, what Neyman’s usage of the term ”confidence” was gesturing toward: a non-oracle observer’s best guess as to how many intervals like the one we have constructed, or the ones we are about to construct, will succeed in covering the parameter. Separating our probabilistic views of confidence procedures in this way—into (i) design-level probabilities under $P_{\theta}$, (ii) degenerate conditionals under $P_{\theta}$ given the full data, and (iii) information-indexed predictive probabilities for the coverage indicator—would ease some of the tension in interpreting constructed intervals ex post, since we could then specify which view we are referring to when we talk about, for example, the “probability” that a single interval has covered the parameter. It would also complement, but not conflict with, views raised by authors who have extended the notion of confidence in other ways, for example, as an extended likelihood [26], or as the basis for CI-based distributional inference [32], as well as those who have rebuilt Fisher’s original notion of fiducial inference [6, 7] with modern methods and philosophical sensibilities [15, 14, 28]. The distinction between this perspective and those is that it treats confidence not as something that might help us learn about value of $\theta$ or its likely distribution, but rather as a predictive probability of the underlying coverage event itself, given that the data, model, and estimator are fixed. In other words, it is simply a model-based guess as to whether the current interval belongs to the subset of intervals the procedure can generate that cover the parameter, not a guess about where, numerically, the parameter actually lies.
To show how such a framing might take shape mathematically, I explore this notion further in a separate paper.

### v3

@@
In Neyman’s original formulation, a $1-\alpha$ confidence interval procedure is justified by its long-run coverage properties, and a single realized interval is often described only by the slogan that it either covers the parameter or it does not. On this view, post-data probability statements about the coverage of an individual interval are taken to be conceptually out of bounds. In this paper, I present two arguments against treating that ”either-or” reading as the only legitimate interpretation of confidence. The first is intuitive, via a set of thought experiments in which the same joint probability model is used to compute both forward-looking and backward-looking probabilities for occurred-but-unobserved events. The second is more formal, recasting the standard confidence-interval construction in terms of infinite sequences of trials and their associated $0/1$ coverage indicators. In that representation, the design-level coverage probability $1-\alpha$ and the degenerate conditional probabilities given the full data appear simply as different conditioning levels within the same model. I argue that a strict behavioristic reading that privileges only the latter is in tension with the very mathematical machinery used to define long-run error rates. I conclude that the familiar “either it covers or it does not” slogan is too restrictive, and that frequentist confidence theory permits a broader class of post-data probability statements than is usually acknowledged.
@@ 1.1 Background
The tension between the accepted interpretation of CIs and the supposedly-fallacious one can be recast more generally as a statement about events we know have occurred, but whose outcomes we have not observed [20]: for frequentists, randomness, and thus probability, lives in the sampling process and not our knowledge of the outcomes, and so once a sample has been drawn, the ex-ante (i.e., ”pre-data”) probability has collapsed to some value in $\{0,1\}$. Although this is mathematically true, it should also give practicing statisticians some pause, because we happily use frequentist methods for statistical inference in exactly this kind of real-world scenario. . Take, for example, the case of medical diagnosis: given that a patient tests positive on, say, a rapid diagnostic test for the influenza virus, what is the probability that she actually has it? If we stay consistent with our interpretation of CIs, we should also say no probability statement may now be made: given her true underlying health state, the patient either does or does not have the flu, and there is no probability left to assign, because all of the randomness in the sampling process has now been exhausted. Following the ”either-or” logic here, though, would substantially limit the clinical value of the diagnostic test in guiding care, and it would obviate the effort epidemiologists and statisticians put in to estimating the test’s positive predictive value ($PPV$) in the first place, neither of which seems particularly desirable.
@@ 1.2 Paper overview and contributions
In what follows, I deliberately adopt a rather strict reading of Neyman’s slogan–in a nutshell, that ex post probabilities of coverage are entirely out of bounds–and treat it as a normative rule (to be fair, though, this is how many, if not most, instructive pieces handle the interpretation, see e.g., [2, 17, 22, 6, 1, 12] for examples). The arguments will form a kind of reductio: if we really insist on this kind of rule, we run into uncomfortable constraints on other frequentist uses of probability. The alternative I suggest, and perhaps this paper’s main contribution, is that we keep Neyman’s (very useful) conception of long-run error control, but that we loosen what we are allowed to say about individual coverage events, especially ex post.
@@ 3.1 Confidence intervals and coverage indicators
with $Z_{i}(\omega)\in\{0,1\}$ for every $i$. From this viewpoint, a single realized world $\omega$ contains a fully-fixed infinite sequence of intervals and coverage outcomes; all of the randomness in the model lives in the draw of $\omega$ itself, which determines every element of the sequence simultaneously. The strong law of large numbers applied to $(Z_{i})_{i\geq 1}$ yields
@@ 4.2 What is confidence, then?
The discussion above suggests that part of the trouble comes from conflating distinct layers of probability that the model makes available. We clearly have at least two model-based layers to work with when interpreting confidence intervals: at the design level, there is the unconditional coverage probability $1-\alpha$ attached to the procedure as a whole, as well as each of its constituent runs; and at the ex post level, conditioned on the truth value of the coverage event for a particular interval, there are the degenerate probabilities 0 and $1$ corresponding to non-coverage and coverage, respectively. Both are properties of the same probability measure $P_{\theta}$, viewed at different conditioning levels. The design-level quantity arguably holds up better under scrutiny, as it avoids the multi collective complications noted above. For some distributions and estimators, though, a bare report of $1-\alpha$ can feel somewhat strange ex post. The “trivial” interval, which returns the real line with probability $1-\alpha$ and the empty set with probability $\alpha$, is perhaps the most familiar example, but Welch’s uniform confidence interval [24] and Basu’s construction based on a $\theta$-free $X$ [3] are also instructive.
@@
In such cases, I would suggest that, while the two model-based layers are certainly still in play—considered only as a random draw, any interval generated by the procedure does, by definition, inherit the design-level coverage probability—it may be more natural to admit a third layer of probability: that of a single-case predictive probability of the coverage event itself, made in light of the information available. From this perspective, our $1-\alpha$ report ex ante can be understood as a default model-based assessment of the coverage indicator for the interval to be constructed, distinct from the interval’s realized coverage status under the model. In some cases, features of both the design and the numerical value of the realized interval may suggest a more refined ex post assessment. In that richer information state, the relevant probability need not equal $1-\alpha$, but it also need not collapse to $\{0,1\}$: it may instead be an intermediate probability statement about the coverage event for the realized interval relative to other intervals generated under the design that are similar with respect to some $\theta$-free feature of the data or interval.
@@
Along these lines, I suggest that this kind of predictive assessment is, in essence, what Neyman’s usage of the term “confidence” was gesturing toward: a non-oracle observer’s best guess as to how many intervals like the one we have constructed, or the ones we are about to construct, will succeed in capturing the parameter. Separating our probabilistic views of confidence procedures in this way—into (i) design-level probabilities under $P_{\theta}$, (ii) degenerate conditionals under $P_{\theta}$ given the full data, and (iii) information-indexed probability statements about the coverage indicator—would ease some of the tension in interpreting constructed intervals ex post, since we could then specify which view we are referring to when we talk about, for example, the “probability” that a single interval has covered the parameter.
To explore this third perspective more fully, including a more explicit mathematical treatment, I take it up in a separate companion paper [13].